OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 5x2 + 20x - 120 = 0

OpenStudy (across):

Listen to me, cherry33. This may be the most important moment of your life. Commit to it.

OpenStudy (anonymous):

I think you meant to say (x+c)^2 right?

OpenStudy (anonymous):

But anyways, first take out a 5

OpenStudy (across):

\[5x^2+20x-120=0,\]\[5x^2+20x=120,\]\[x^2+4x+4=24+4,\]\[(x+2)^2=28.\]

OpenStudy (anonymous):

yea i keep tryin do it but i keep gettin an answer that isnt there im really bad at these